If a salesman's average is a new order every other week, he will
break the office record of the year. However, after 28 weeks, he is six
orders behind schedule. In what proportion of the remaining weeks does
he have to obtain a new order to break the record?

Would you please answer this question as soon as possible?i have an exam tomorrow.

Answer by khwang(438)

Assume there are 52 weeks in one year.
Since he supposed to have a new order for every two weeks, he
needs 52/2 = 26 orders to break the office record.

Now after 28 weeks,he has got only 28/2 -6 = 8 orders.
Hence,he needs 26-8 = 18 new orders in the remaining 24= 52-28 weeks to break the office record.

Compute 24 orders/18 weeks = 4/3 orders/week ,
we see that averagely he has a new order for
every 4/3 weeks in the remaining weeks to break the office record.
Kenny